### v1

# SMITE: Segment Me In TimE

Abstract
Segmenting an object in a video presents significant challenges. Each pixel must be accurately labeled, and these labels must remain consistent across frames. The difficulty increases when the segmentation is with arbitrary granularity, meaning the number of segments can vary arbitrarily, and masks are defined based on only one or a few sample images. In this paper, we address this issue by employing a pre-trained text to image diffusion model supplemented with an additional tracking mechanism. We demonstrate that our approach can effectively manage various segmentation scenarios and outperforms state-of-the-art alternatives. The project page is available at https://segment-me-in-time.github.io/.
Amirhossein Alimohammadi1, Sauradip Nag1, Saeid Asgari Taghanaki1,2,
Andrea Tagliasacchi1,3,4, Ghassan Hamarneh1, Ali Mahdavi Amiri1
1Simon Fraser University 2Autodesk Research 3University of Toronto 4Google DeepMind
Figure 1: SMITE. Using only one or few segmentation references with fine granularity (left), our method learns to segment different unseen videos respecting the segmentation references.
1 Introduction
Segmenting an object in a video poses a significant challenge in computer vision and graphics, frequently employed in applications such as visual effects, surveillance, and autonomous driving. However, segmentation is inherently complex due to variations in a single object (scale, deformations, etc.), within the object class (shape, appearance), as well as imaging (lighting, viewpoint). In addition, difficulty arises due to the segmentation requirements, such as its granularity (i.e., number of segments), as demanded by the downstream tasks. For example, in face segmentation, one VFX application might need to isolate the forehead for wrinkle removal, while another, such as head tracking, might treat it as part of the whole face. Creating a comprehensive dataset for every possible segmentation scenario to develop a supervised segmentation technique is extremely time-consuming and labor-intensive. Therefore, there is a need to segment images or videos based on a reference image. We call this type of segmentation flexible granularity.
When flexible granularity segmentation is applied on a large scale, it significantly improves downstream tasks such as VFX production, which involves managing numerous shots and videos. By segmenting one or a few reference images only once and then using those images to segment any video that features a target object of the same class, we can eliminate the need for separate segmentation of each video, thereby making the process far more efficient. In this paper, we tackle the challenge of video segmentation using one or few reference images that are not derived from the video frames themselves. For example, as shown in Fig. 1, a few annotated images are provided as references to our model, and our method, SMITE, successfully segments videos, exhibiting an object from the same class, in the same level of granularity. Importantly, none of the frames from these videos are included in the reference images, yet SMITE is capable of segmenting the videos with objects that exhibit different colors, poses, and even occlusions. This is an important feature when working with large-scale videos requiring consistent segmentation (such as VFX videos needing the same enhancements), since there is no need for manual intervention to segment each video’s frames.
While recent work has explored flexible granularity segmentation for objects in images by leveraging the semantic knowledge of pretrained text-to-image diffusion models Khani et al. (2024), the complexity increases in videos. Ensuring label consistency across frames and managing instances where image segmentation may fail to produce accurate results require additional considerations.
To tackle these challenges and achieve consistent segmentation across frames, we utilize the semantic knowledge of pretrained text-to-image diffusion models, equipped with additional temporal attentions to promote temporal consistency. We also propose a temporal voting mechanism by tracking and projecting pixels over attention maps to maintain label consistency for each pixel. This approach results in segmentations with significantly reduced flickering and noise compared to per-frame segmentation methods while segments still follow the reference images thanks to our low-pass regularization technique that ensure preserving the structure of segments provided by attention maps and optimized according to the reference images.
Moreover, rather than simply optimizing a token for each segment Khani et al. (2024), we also fine-tune cross-attentions to enhance segmentation accuracy and better align with the reference images. Consequently, our method not only supports videos with temporal consistency but also outperforms flexible granularity image segmentation techniques in segmenting a single image.
We validate our design choices and methodology through comprehensive experiments detailed in the paper. As existing datasets with arbitrary semantic granularity are lacking, we introduce a small dataset, SMITE-50, to demonstrate the superior performance of our method against baselines. Additionally, we conduct user studies that highlight our method’s effectiveness in terms of segmentation accuracy and temporal consistency.
2 Related Work
Part-based semantic segmentation.
In computer vision, semantic segmentation, wherein a class label is assigned to each pixel in an image, is an important task with several applications such as scene parsing, autonomous systems, medical imaging, image editing, environmental monitoring, and video analysis (Sohail et al., 2022; He et al., 2016; Chen et al., 2017a; Zhao et al., 2017; He et al., 2017; Chen et al., 2017b; Sandler et al., 2018; Chen et al., 2018; Ravi et al., 2024). A more fine-grained derivative of semantic segmentation is semantic part segmentation, which endeavors to delineate individual components of objects rather than segmenting the entirety of objects. Despite notable advancements in this domain (Li et al., 2023; 2022), a limitation of such methodologies is their reliance on manually curated information specific to the object whose parts they aim to segment. To solve the annotation problem, some works(Pan et al., 2023; Wei et al., 2024) proposed open-set part segmentation frameworks, achieving category-agnostic part segmentation by disregarding part category labels during training. Building on this, further works such as SAM (Kirillov et al., 2023), Grounding-SAM (Ren et al., 2024) explored utilizing foundation models to assist in open-vocabulary part segmentation. However, most of these methods can only segment the parts that are semantically described by text. With the influx of Stable Diffusion (SD) based generative segmentation approaches (Khani et al., 2024; Namekata et al., 2024), such issues have been partly solved by allowing SD features to segment semantic parts at any level of detail, even if they cannot be described by text. Despite such progress, applying such fine-grained segmentations on videos is challenging and unexplored. Our proposed SMITE presents the first part-segmentations in videos wherein it segments utilizing the part features from a pre-trained SD and generalizes it to any-in-the wild videos.
Video segmentation. Video segmentation methods can be categorized as video semantic segmentation (VSS) (Zhu et al., 2024; Zhang et al., 2023a; b; Li et al., 2024; Ke et al., 2023; Wang et al., 2024), video instance segmentation (VIS) (Yang et al., 2019) and video object segmentation (VOS) (Xie et al., 2021; Wang et al., 2021b; Cheng et al., 2021a; b; Bekuzarov et al., 2023). VSS and VIS extends image segmentation to videos, assigning pixel labels across frames while maintaining temporal consistency despite object deformations and camera motion. VOS, in contrast, focuses on tracking and isolating specific objects throughout the video. Both tasks leverage temporal correlations through techniques like temporal attention (Mao et al., 2021; Wang et al., 2021a), optical flow (Xie et al., 2021; Zhu et al., 2017), and spatio-temporal memory (Wang et al., 2021b; Cheng & Schwing, 2022). Recent efforts, such as UniVS (Li et al., 2024), proposes unified models for various segmentation tasks, utilizing prior frame features as visual prompts. However, these methods struggle with fine-grained part segmentation and generalization to unseen datasets (Zhang et al., 2023b). Bekuzarov et al. (2023) leverage a spatio-temporal memory module and a frame selection mechanism to achieve high-quality video part segmentation with partial annotations. However, it requires frame annotations from the same video complicating video segmentation at scale. In contrast, we only need segmentation references for a few arbitrary selected images and we can segment an unseen given video respecting the segementation references. Therefore, per video manual annotation is not needed in our method.
Video diffusion models. Recently, diffusion models (Ho et al., 2020; Song et al., 2020a; b) have gained popularity due to their training stability and have been used in various text-to-image (T2I) methods (Ramesh et al., 2021; 2022; Saharia et al., 2022; Balaji et al., 2022), achieving impressive results. Video generation (Le Moing et al., 2021; Ge et al., 2022; Chen et al., 2023b; Cong et al., 2023; Yu et al., 2023; Luo et al., 2023) can be viewed as an extension of image generation with an additional temporal dimension. Recent video generation models (Singer et al., 2022; Zhou et al., 2022; Ge et al., 2023; Nag et al., 2023; Cong et al., 2024) attempt to extend successful text-to-image generation models into the spatio-temporal domain by inflating the T2I UNet. VDM (Ho et al., 2022) adopted this inflated UNet for denoising while LDM (Blattmann et al., 2023) implement video diffusion models in the latent space. Video diffusion models can be categorized into inversion-based and inversion-free methods. Inversion-based approaches (Cong et al., 2024; Jeong & Ye, 2023) use DDIM inversion to control attention features ensuring temporal consistency, while inversion-free methods (Zhang et al., 2023c) focus on more flexible conditioning, wider compatibility and better generation quality. However, inversion-free methods can suffer from flickering due to a lack of DDIM inversion guidance. Recent works (Wang et al., 2024; Zhu et al., 2024) explored inversion-free T2V diffusion models to segment objects in videos, but they fail to generate fine-grained segments and often produce flickering segmentation results. Building on this, our method solves the seminal problem of video part-segmentation which combines an inversion-free model coupled with point-tracking algorithms (Karaev et al., 2023) to generate consistent, generalizable, flicker-free segmentations.
3 Preliminaries
Latent diffusion models and WAS maps. Latent Diffusion Models perform the denoising operation on the latent space of a pretrained image autoencoder. Each latent pixel corresponds to a patch in the generated image. Starting from pure random noise $z_{T}$, at each timestep $t$, the current noisy latent $z_{t}$ is passed through a denoising UNet $\epsilon_{\theta}$, which is trained to predict the current noise $\epsilon_{\theta}(z_{t},y,t)$ using text prompt $y$. In each block, the UNet employs residual convolution layers to generate intermediate features, which are then fed into attention layers. These attention layers average various values based on pixel-specific weights. The cross-attention layers (denoted by $A_{ca}$) incorporate semantic contexts from the prompt encoding, whereas the self-attention layers (denoted by $A_{sa}$) leverage global information from the latent representation itself. SLiMe (Khani et al., 2024) demonstrated that text embeddings can be learned from a few images to segment other unseen images by leveraging both cross attention and self-attention layers in a novel representation called Weighted Accumulated Self-Attention (WAS) map, $S_{\text{WAS}}$. This is defined as follows:
$S_{\text{WAS}}=\text{Sum}(\text{Flatten}(R_{ca})\odot A_{sa}),$ (1)
where $R_{ca}$ is the downsampled latent of $A_{ca}$.
Inflated UNet. A T2I diffusion model, such as LDM (Rombach et al., 2022), usually utilizes a U-Net(Ronneberger et al., 2015) architecture, which involves a downsampling phase, followed by an upsampling with skip connections. The architecture consists of layered 2D convolutional residual blocks, spatial attention blocks, and cross-attention blocks that incorporate textual prompt embeddings. To extend the T2I model for T2V tasks, the convolutional residual blocks and spatial attention blocks are inflated. Following earlier approaches (Cong et al., 2024; Wu et al., 2022), the $3\times 3$ convolution kernels in the residual blocks are adjusted to $1\times 3\times 3$ by introducing a pseudo temporal channel. To enhance the temporal coherence, we further extend the spatial self-attention mechanism to the spatio-temporal domain. The original spatial self-attention method focused on patches within a single frame. However, in inflated UNet, we use all patch embeddings from the entire video as the queries, keys, and values. This allows for a full understanding of the video’s context. Additionally, we reuse the parameters from the original spatial attention blocks in the new dense spatio-temporal attention blocks.
4 Method
Here, we first introduce and formalize our method (SMITE) designed for achieving temporally consistent video segmentation with varying levels of granularity, guided by one or more reference images (Sec. 4.1). To achieve this and capture fine-grained segments, we first propose a new training strategy applied on the inflated UNet (Sec. 4.2). The segmentation obtained from the inflated UNet may lack temporal consistency. To address this, we employ a voting mechanism guided by a tracking method that is projected onto the attention maps. However, relying solely on tracking to adjust the segments might lead to deviations from the provided samples. To mitigate this, we incorporate a frequency-based regularization technique to maintain detailed segmentations across frames (Sec. 4.3). Since tracking and frequency-based regularization may pull the segmentation in different directions, we use a energy based guidance optimization technique to balance both approaches (Sec. 4.4).
4.1 Problem Setting
Problem statement. Given one or few images, $\mathcal{I}_{n}\in\mathbb{R}^{H\times W\times 3}$, of a subject, along with its segment annotations $\mathcal{Y}_{n}=\{\mathcal{Y}^{i}_{n}|i=1:K\}$ where $\mathcal{Y}^{i}$ denotes binary segment mask $i$, and $n$ and $K$ are respectively the number of images and segments. Our objective is to learn temporally consistent segments of the subject for a given video $\mathcal{V}=\{v^{M}_{j=1}\}$ where $v_{j}$ represents video frames.
Figure 2: SMITE pipeline. During inference (a), we invert a given video into a noisy latent by iteratively adding noise. We then use an inflated U-Net denoiser (b) along with the trained text embedding as input to denoise the segments. A tracking module ensures that the generated segments are spatially and temporally consistent via spatio-temporal guidance. The video latent $z_{t}$ is updated by a tracking energy $\mathcal{E}_{track}$ (c) that makes the segments temporally consistent and also a low-frequency regularizer (d) $\mathcal{E}_{reg}$ which guides the model towards better spatial consistency.
Our framework. We use Stable Diffusion’s (SD) semantic knowledge to learn the segments defined by few images and then generalize them to the video of the subject in any pose, color or size. This implies that the model needs to share information across frames to enforce temporal consistency. Differently from the UNet structure used in Khani et al. (2024), we apply an inflation to the T2I models across the temporal dimension (Wu et al., 2022) to enable temporal attention across all video frames. Also, we incorporate a tracking module combined with low-frequency regularization to enhance spatio-temporal consistency across frames. The overall inference pipeline of our model, SMITE, is illustrated in Fig. 2.
4.2 Learning generalizable segments
We first learn segments provided by the reference images of the subject by optimizing a text embedding for each segment of the reference images that can be used for segmenting the given videos. We also fine tune the cross attentions of the SD to better match the provided segments since the text embeddings alone may not be able to fully capture the masks’ details.
Learning text embeddings. We begin by passing the reference images $\mathcal{I}$ and text embeddings $\mathcal{T}$ into into SMITE (denoted by $\psi(.)$) . Similar to SLiMe (Khani et al., 2024), we obtain image resolution WAS maps (denoted by $\mathcal{S}$) from our inflated UNet as follows:
$\mathcal{S}=\psi_{\theta}(\mathcal{I},\mathcal{T})\hskip 7.22743pt\text{,}\hskip 7.22743pt\hat{Y}=argmax(\mathcal{S}),$ (2)
[…]
Network fine-tuning. Only learning text embeddings fails to capture complex granularities. To extract more customizable fine-grained segments from the video, we need to fine-tune the existing SD weights (denoted by $\theta$) using the available provided segment annotations. As shown in the inset figure, optimizing solely the text embeddings struggles with asymmetrical segmentation (e.g., segmenting only one eye). We hypothesize that such issues arises because the model may get stuck in local minima when relying exclusively on text embeddings for optimization. To mitigate this, we update the cross-attention layers $A_{ca}$ in the UNet along with the text-embedding but do so in two phases. First, the model is frozen, while the text embeddings corresponding to the segments denoted by $\mathcal{T}$ are optimized using a high learning rate. Thus, an initial embedding ($\mathcal{T}^{*}$) is achieved quickly without detracting from the generality of the model, which then serves as a good starting point for the next phase. Second, we unfreeze the cross-attention weights and optimize them along with the text embeddings, using a significantly lower learning rate. This gentle fine-tuning of the cross-attention and the embeddings enables faithful generation of segmentation masks with varied granularity. We use the same combination of losses ($\mathcal{L}_{\text{CE}}$, $\mathcal{L}_{\text{MSE}}$ and $\mathcal{L}_{\text{LDM}}$) used in SLiMe (Khani et al., 2024) for network finetuning. This results in an optimized SMITE model (denoted by $\Psi_{\theta}^{*}$) which can generate segments that is generalizable across different videos.
4.3 Temporal Consistency
To enhance temporal consistency, we use temporal attention in SMITE’s Inflated UNet denoiser. While it improves temporal consistency compared to independent frame processing (Fig. 3(a)), inconsistencies remain due to the need for segmenting at different granularities, often with imprecise boundaries. These inconsistencies can cause flickering or unnatural transitions (Fig. 3(b)).
[…]
[height=2.3in,width=0.45loop, autoplay]10img/ablation_gif/00039
Figure 3: Best viewed in Adobe Acrobat.
Figure 4: Segment tracking module ensures that segments are consistent across time. It uses co-tracker to track each point of the object’s segment (here it is nose) and then finds point correspondence of this segment (denoted by blue dots) across timesteps. When the tracked point is of a different class (e.g,. face) then it is recovered by using temporal voting. The misclassified pixel is then replaced by the average of the neighbouring pixels of adjacent frames. This results are temporally consistent segments without visible flickers.
Segment tracking and voting. The first step to ensure segment consistency is tracking the segments across time. Point tracking methods like CoTracker (Karaev et al., 2023) are well-suited for our approach because they use point correspondences to minimize pixel drift over time. However, since our segments come from attention maps, tracking needs to occur directly on these maps. Since CoTracker is trained on spatial domains, we first apply it on the frames of video $\mathcal{V}$ using a sliding window, project the results onto the attention maps via simple scaling, and then update the attention maps accordingly. A pixel’s label in the attention map is updated based on the most frequent label it receives across the visible frames in the window. Formally, for a pixel with coordinates $(x_{t},y_{t})$ on the frame latent at timestep $t$, its coordinates on all subsequent frame latents in the window can be derived from the tracker. The coordinates are linked, and the trajectory sequence can be presented as:
$\{(x_{t-\frac{w}{2}},y_{t-\frac{w}{2}}),...,(x_{t},y_{t}),...,(x_{t+\frac{w}{2}},y_{t+\frac{w}{2}})\}=\phi(\mathcal{P}(\mathcal{V})),\hskip 7.22743pt$ (3)
where $\mathcal{P}$ is the tracker, $\phi(.)$ is the scaling operator, and $w$ denotes the length of the sliding window. We choose the point corresponding to the center frame of the window $(x_{t},y_{t})$ as the first point and track bidirectional correspondences within the window. Each of these trajectory points $(x,y)$ is assigned a segment label $\hat{Y}$. However, certain pixels in the window may disappear and reappear over time as shown in Fig 4. To handle such scenarios, we discard labels for which the tracking pixel is invisible and then use temporal voting to update its segment labels. Therefore, we update the WAS maps (denoted by $\mathcal{S}_{\text{WAS}}$) corresponding to the tracking pixel $(x_{t},y_{t})$ and use the following formula to assign the labels:
$\mathcal{S}_{\text{Tracked}}(x_{t},y_{t})=\text{Avg}(\mathcal{S}(x_{l},y_{l})\mid\hat{Y}(x_{l},y_{l})=F)\hskip 7.22743pt\forall l\in(t-\frac{w}{2},t+\frac{w}{2}),$ (4)
where $F$ denotes the most frequent label in the window. After correcting the unstable pixel tracking over multiple-step denoising, we obtain a consistent set of segments ($\mathcal{S}_{tracked}$). Note that the window size $w$ is small (e.g, seven), hence to track the segments for the entire video, we need to slide the windows and track in an iterative fashion. More details are provided in the supplementary material.
Intuitively, if we would have relied only on cross-attention for segmentation, we could have adjusted it directly to align with the tracking output. However, since segments $\mathcal{S}$ are derived from the combination of cross and self-attention, direct manipulation is challenging. To tackle this, we propose an energy function to measure how well the segments before tracking (denoted by $\mathcal{S}$) align with the segments after correcting (denoted by $\mathcal{S}_{tracked}$) it using temporal voting. This energy is defined as:
$\mathcal{E}_{\text{Tracking}}=\textit{CE}(\text{S},\text{S}_{\text{tracked}}),$ (5)
where $CE(.)$ represents cross-entropy objective. This optimization ensures that the segment drift is corrected throughout the video to make it temporally consistent with low flicker.
Low-pass regularization. When tracking is applied, label modifications should be repeated through several denoising steps to ensure full temporal label propagation and achieve better consistency. This process may cause deviations (Fig. 3) from the original segmentation provided by the reference images that is captured in the segments ($\mathcal{S}$) in the initial denoising step. Our goal is to preserve the overall structure of the segments $\mathcal{S}$ from the WAS maps while smoothing boundary transitions for temporal consistency via tracking. Low Pass Filters (LPF) have been effective in video generation to maintain temporal correlations and spatial frame structure (Wu et al., 2023; Si et al., 2024). Therefore, we use an LPF on the finally denoised segments $\mathcal{S}$ ensuring the structure of the LPF on the segments $\mathcal{S}$ predicted in the initial denoising step (denoted by $\mathcal{S}_{ref}$) is respected by the following function:
$\mathcal{E}_{\text{Reg}}=\left\|\omega(\mathcal{S})-\omega(\mathcal{S}_{ref})\right\|_{1}$ (6)
where $\omega(S)=\mathcal{H}_{l}\odot\text{DCT}_{3D}(\text{S})$, $\text{DCT}_{3D}$ refers to Discrete Cosine Transform and $\mathcal{H}_{l}$ refers to an LPF. This ensures that the segments structure is progressively corrected across the video frames at the end of the denoising phase as illustrated in Fig. 3.
4.4 Spatio-Temporal Guidance
Our overall energy function is minimized by backpropagating through the diffusion process, as described in Chen et al. (2023a); Safaee et al. (2023), which updates the latent representation to achieve more consistent segmentation over time:
$\mathcal{E}_{\text{Total}}=\lambda_{\text{Tracking}}\cdot\mathcal{E}_{\text{Tracking}}+\lambda_{\text{Reg}}\cdot\mathcal{E}_{\text{Reg}},$ (7)
where $\lambda_{\text{Tracking}}$ and $\lambda_{\text{Reg}}$ are the coefficients for the tracking and reference loss functions, respectively. Optimizing this function encourages better spatial and temporal consistency in the WAS attention maps (S). Specifically, we update the latent $z_{t}$ using gradient descent on $\mathcal{E}_{\text{Total}}$:
$z_{t}^{\prime}\leftarrow z_{t}-\alpha_{t}\cdot\nabla_{z_{t}}\mathcal{E}_{Total},$ (8)
where $\alpha_{t}$ is the learning rate. Results indicate that this strategy leads to superior performance compared to when we either omit the voting mechanism or do not apply low-pass regularization.
5 Results and Experiments
Dataset and benchmark. To evaluate our method, we introduce a benchmark dataset called SMITE-50, primarily sourced from Pexels. SMITE-50 features multi-granularity annotations and includes visually challenging scenarios such as pose changes and occlusions. To our knowledge, no existing datasets focus exclusively on multi-granular and multi-segment annotations. While the PumaVOS dataset Bekuzarov et al. (2023) contains a limited number of multi-part annotated videos across a diverse range, it does not offer multiple videos for specific granularities and categories.
Figure 5: SMITE-50 Dataset sample.
We focus on three main categories: (a) Horses, (b) Human Faces, and (c) Cars, encompassing 41 videos. Each subset includes ten segmented reference images for training and densely annotated videos for testing. The granularity varies from human eyes to animal heads, etc. relevant for various applications such as VFX (see Fig. 5). All segments are labeled consistently with the part names used in existing datasets. Additionally, we provide nine challenging videos featuring faces with segments that cannot be described textually, as shown in Fig. 3 (Non-Text). Overall, our dataset comprises 50 video clips, each at least five seconds long. For dense annotations, we followed a similar approach to (Ding et al., 2023; Bekuzarov et al., 2023), creating masks for every fifth frame with an average of six parts per frame across three granularity types (more info in Appendix). While PumaVOS dataset Bekuzarov et al. (2023) has 8% annotations, our SMITE-50 dataset has 20% dense annotations.
Evaluation protocol. In our setting, few reference images per class are used to train SMITE and then it is evaluated on videos from the same categories but not the same objects in the training data. For all the cases, we report the standard metrics (higher is better): Mean Intersection Over Union (mIOU), and Contour Accuracy $F_{measure}$ respectively.
Figure 6: Visual comparisons with other methods demonstrate that SMITE maintains better motion consistency of segments and delivers cleaner, more accurate segmentations. Both GSAM2 and Baseline-I struggle to accurately capture the horse’s mane, and GSAM2 misses one leg (Left), whereas our method yields more precise results. Additionally, both alternative techniques create artifacts around the chin (Right), while SMITE produces a cleaner segmentation.
Quantitative comparison. Since there are no available few-shot video part segmentation methods that can be applied to our setting, we build the following two state-of-the-art baselines for quantitative comparison: (1) Few-shot Image segmentation based approach, in which SLiMe (Khani et al., 2024) is applied to the video frame-by-frame. We term this as Baseline-I. (2) Grounded SAM2 (Ren et al., 2024; Ravi et al., 2024), a recently introduced zero-shot foundation model for video segmentation approach which is applied directly on the video. For Baseline-I, we applied the same training procedure and used SMITE-50 dataset. As apparent in Tab. 1, our method produces the most accurate results in comparison with other methods across all categories and all metrics.
Generally, XMem++ (Bekuzarov et al., 2023) cannot be utilized in our setting since we do not want to use any frames from the provided video. However, in the Appendix, we included a comparison with XMem++ on a subset of their proposed dataset PUMaVOS with flexible granularity to provide a contrast and show that our method is effective on datasets beyond SMITE-50. Although XMem++ is a semi-supervised technique, our method performs comparably and even outperforms it with fewer frames (e.g., a single shot). When XMem++ is given more frames per video (e.g., 10), its performance slightly surpasses ours, which is expected since SMITE does not require any substantial video pre-training. However, with a smaller number of frames (e.g., one or five), our method either outperforms or matches the performance of XMem++. Additionally, we report our performance on the image segmentation task in the Appendix, showing that we outperform SLiME (Baseline-I), highlighting the effectiveness of our design choices, including the cross attention optimization.
Table 1: Quantitative evaluation on SMITE-50 dataset. The results are presented for each category (Face, Horse, Car, Non-Text) having 10 reference image during training.
Faces Horses Cars Non-Text
Methods $\textbf{F}_{meas.}$ mIOU $\textbf{F}_{meas.}$ mIOU $\textbf{F}_{meas.}$ mIOU $\textbf{F}_{meas.}$ mIOU
Baseline-I 0.81 72.95 0.64 65.48 0.57 61.38 0.67 66.69
GSAM2 0.73 63.28 0.76 72.76 0.64 63.56 - -
Ours 0.89 77.28 0.79 75.09 0.82 75.10 0.77 73.08
Qualitative comparisons. Qualitative comparisons are presented in Fig. 6. GSAM2 struggles to locate the boundaries accurately and produces coarse segments. SLiMe can usually segment the first frame accurately but struggles to preserve temporal consistency. Our method produces the best segmentation maps in terms of segmentation quality and temporal consistency. The maps have sharper boundaries and clean clusters compared to other methods. Our SMITE performs better than its SLiMe counterpart. In Fig. 7, we present results from other categories, demonstrating the versatility of our model in handling various object categories. As shown, despite significant differences in pose, expression, gender, and other properties between the video and the annotated images, our method still delivers high-quality results. This is particularly evident in the pineapple example, where the object is cut in half, yet the method successfully tracks it and produces accurate segmentation. For more challenging cases, we include results from additional difficult videos in Fig. 8. In one instance, the ice cream cone is occluded by a paper napkin, and the ice cream itself is obscured and blended by a face, yet SMITE is still able to generate correct results. Furthermore, the turtle nearly blends with the background in terms of color and visual patterns, but our method successfully tracks the segments.
Ablation study. To demonstrate the effectiveness of our design choices, we conducted three ablation studies. The first ablation on the Car dataset (Tab. 3), illustrates that combining $\mathcal{E}_{\text{Tracking}}$ and $\mathcal{E}_{\text{Reg}}$ improves the results. It’s important to note that tracking without low pass regularization leads to unsatisfactory outcomes, as results tend to align too closely with the voting method. Tab. 3 shows the impact of cross attention tuning also on the Car dataset. By optimizing both cross attention and text embedding, we achieve the highest accuracy, demonstrating the effectiveness of our tuning strategy. We have also ablated the number of images for training our few shot setting $k=1/5/10$. While our performance improves with more training images, our one shot setting still performs reasonably well (Tab. 5).
Table 2: Loss ablation. When two losses are combined, the best performance is achieved.
$\mathcal{E}_{\text{Tracking}}$ $\mathcal{E}_{\text{Reg}}$ mIOU
69.85
75.10
Table 3: Cross attention optimization improves the quality of segmentation.
Finetuning Strategy mIOU
Text Embedding 74.00
Cross Attention + Text 75.10
User study. We conduct a user study because human judgment is best for assessing perception when both temporal consistency and faithfulness to segmentations are important. We collected 16 segmented videos (4 from Non-Text category) and asked 40 participants to rank the methods (i.e., 1 best and 3 worst) based on segmentation quality (i.e., fidelity to the segmentation reference) and motion consistency (i.e., reduced flicker), encouraging them to prioritize segmentation quality in their evaluations in terms of a tie in motion consistency. We do not report the scores of GSAM2 on Non-Text segments as the segments in the videos are not describable by texts. Tab. 5 demonstrates that SMITE achieves higher preference both for textual and non-textual segmentation.
Table 4: Few shot ablation on cars in SMITE-50. The performance increases with more training images but still performs well in one shot setting.
Training sample # mIOU
1-shot 63.03
5-shot 71.55
10-shot 75.10
Table 5: User study. We are ranked the best for both textual and Non-Text classes.
Methods Motion Consistency
Horse, Car, Face Non-Text
Baseline-1 2.58 2.37
GSAM2 2.13 -
Ours 1.19 1.10
6 Conclusion
In this work, we introduce SMITE, a video segmentation technique that supports flexible granularity segmentation of objects within a video. SMITE leverages the semantic knowledge of a pre-trained text-to-image diffusion model to segment a video with minimal additional training (i.e., only a few images), while utilizing the temporal consistency of the video to maintain motion consistency within the segments. Notably, SMITE can be trained with a few images that are not necessarily from the video, yet it effectively segments objects within an unseen video. To better show the capabilities of our method, we also collected a flexible granularity dataset called SMITE-50 that will be publicly available along with our code. Through various quantitative and qualitative experiments, as well as user studies, we demonstrate the effectiveness of our method and its components. However, our method still faces limitations that motivate future work. Specifically, SMITE does not perform well when the target objects or segments are too small, and its performance degrades when dealing with with low video resolution. Addressing these issues presents interesting avenues for future research. Additionally, while we use Co-tracker to track segments, exploring other tracking systems and evaluating their performance would be valuable.
Figure 7: Additional results. We visualize the generalization capability of SMITE model (trained on the reference images) in various challenging poses, shape, and even in cut-shapes.
Figure 8: Segmentation results in challenging scenarios . SMITE accurately segments out the objects under occlusion (”ice-cream") or camouflage (”turtle") highlighting the robustness of our segmentation technique.